# Photon rings and shadows of Kerr black holes immersed in a swirling universe

### v1

@@ 3.1 Light rings
Figure 6: Location of photon rings in the $(r,\theta)$-plane for $a=0.5M$ (upper row) and $a=0.9M$ (lower row). The ergoregions are highlighted in light blue and the vertical dashed line represents the radius of the event horizon. The direction of rotation of the light rings is indicated by their colour : prograde orbits are represented by blue dots, while retrograde ones are represented by red dots. Note that prograde and retrograde here are with respect to the direction of the black hole rotation.
@@ 5 Conclusions
Kerr black holes in a swirling universe are closed form solutions in General Relativity, obtained by well-known transformations from the Kerr black hole seed solution [4]. Unlike the Schwarzschild black holes in a swirling universe, however, the interaction between the black hole spin and the swirling parameter does not yield a (odd) $\mathbb{Z}_{2}$ symmetry of the spacetime. Thus the upper and lower hemispheres are not related by a simple symmetry transformation.
@@
Here we have investigated some properties of this spacetime. First we have recalled the conical singularities, which diverge at a critical value of the modulus of the parameter combination $ajM=0.25$, where both spin parameters enter. Next we have addressed the ergoregions. These also exhibit a notable change in the vicinity of the critical value. Whereas for small $ajM$, upper and lower hemispheres possess disconnected counter-rotating non-compact ergoregions in addition to the common ergoregion of the Kerr black hole, this changes in the vicinity of the critical value. For larger values of $ajM$, the three disconnected regions merge to become one infinitely extended ergoregion with hence only one sense of direction.
Then we addressed the first of our main objectives, the light rings in these KBHSU spacetimes. For the SBHSU spacetime the degenerate Schwarzschild LRs split up in a symmetrical way with increasing swirling parameter and move towards the horizon and the poles. For the KBHSH spacetime the picture is far more interesting. In particular, at about the critical value of $ajM=0.25$ a third LR appears from the axis. But the three LRs co-exist then only in a small range of the spin parameters, until one of the two original LRs disappears. Moreover, in contrast to the SBHSU spacetime, the LRs typically possess different radii.
@@
Finally, we investigated the shadow of KBHSU spacetimes, our second main objective. Since the KBHSU metric is Petrov type I and hence the equations describing geodesic motion are not separable, we had to employ fully numerical methods to obtain the shadows seen by observers outside the black hole. The SBHSU shadows exhibit an odd-$\mathbb{Z}_{2}$ symmetry for observers in the equatorial plane. The interaction between the spins of the KBHSU spacetime does not allow for such odd-$Z_{2}$ symmetric shadows. All shadows are twisted, though, which would make the appearance of such a black hole most unusual.
@@
While the spacetime may not appear very realistic given the well-known features of our universe, it may still have features that might be of relevance. Here we would like to point out that the universe is threaded by huge structures, cosmic filaments, that rotate and thus might be modelled in part by the swirling universe solutions. Moreover, Astorino et al. [4] considered the swirling universe solutions as possibly relevant in the description of the collision of rotating galaxies. Finally, we would like to mention that it has recently been proposed that a rotating universe might solve the Hubble tension [26].
@@
The KBHSU is also interesting from another point of view : it is an example of a space-time in which test particle motion is chaotic [27]. It would hence be an ideal system to quantify and qualify chaotic behaviour. This could lead to a better understanding of the underlying connection between chaotic test particle motion and the algebraic classification of solutions generated via suitable transformations in the Ernst formalism.
@@
Acknowledgments: R.C. would like to thank CAPES for financial support under Grant No: 88887.371717/2019-00.

### v2

@@ 3.1 Light rings
Figure 6: Location of photon rings in the $(r,\theta)$-plane for $a=0.5M$ (upper row) and $a=0.9M$ (lower row). The ergoregions are highlighted in light blue and the vertical dashed line represents the radius of the event horizon. The direction of rotation of the light rings is indicated by their markers : prograde orbits are represented by dots, while retrograde ones are represented by squares. The colors indicate that the LR were obtained from the $H_{\pm}$ potentials. Note that prograde and retrograde here are with respect to the direction of the black hole rotation.
3.2 Light points
An interesting phenomenon can be seen taking place in Fig. 6: the outermost light ring (depicted in orange) changes the sign of its angular velocity. For small values of $jM^{2}$ it is counter-rotating with the black hole, but when the swirling parameter is increased this light ring is captured by the ergoregion and it starts co-rotating with the black hole. Hence, its angular velocity must vanish at some point, indicating the presence of a light point [14]. Such peculiar orbits arise when a light ring sits at the boundary of the ergoregion, there the frame dragging can only be compensated by a particle moving at the speed of light. It is precisely this that happens with a photon at a light point - it is counter-rotating at the speed of light but its speed is balanced by the frame-dragging such that it appears to be at rest when seen by an asymptotic observer.
It is possible to show that the light point occurs precisely at the point where the two disconnected ergoregions merge. To prove this, it is important to note that the light point is a saddle point of $g_{tt}$, hence it satisfies:
$g_{tt}=0\quad\wedge\quad\nabla g_{tt}=0\,.$ (17)
The first condition imposes some constraints on the effective potentials, as:
$H_{\pm}\overset{g_{tt}=0}{=}\frac{-g_{t\varphi}\pm\left|g_{t\varphi}\right|}{g_{\varphi\varphi}}\,$, (18)
so if $g_{t\varphi}>0$ one has $H_{+}=0\,,H_{-}=-2\left|g_{t\varphi}\right|/g_{\varphi\varphi}$, if $g_{t\varphi}<0$ the role of the potentials is reversed. For concreteness, it is assumed that $g_{t\varphi}>0$. From the definition of the effective potentials it also follows that they are related to $g_{tt}$ by:
$g_{tt}=g_{\varphi\varphi}H_{+}H_{-}\,.$ (19)
Hence, its gradient can be written as:
$\nabla g_{tt}=H_{+}\nabla\left(g_{\varphi\varphi}H_{-}\right)+g_{\varphi\varphi}H_{-}\nabla H_{+}\,.$ (20)
As seen before, the ergoregion condition implies that $H_{+}=0$, moreover $H_{-}<0$ and $g_{\varphi\varphi}$ is also positive everywhere outside the axis. And the axis has no common points with the ergosphere so it is impossible to have $g_{tt}=0$ and $g_{\varphi\varphi}=0$ simultaneously. Therefore, to satisfy $\nabla g_{tt}=0$ one must have $\nabla H_{+}=0$, so in summary
[…]
But the conditions on the right hand side are precisely the LR conditions, indicating the presence of a light point ($H_{\pm}=0$) precisely where the ergorregion has a saddle point. Thus, this not only proves that when a light ring coincides with a saddle point of the boundary of the ergoregion it is a light point, but rather that whenever the boundary of the ergoregion has a saddle point, that point corresponds to a light point.
The explicit form of the equations in Eq. (17) is very convoluted and it was not possible to obtain a closed-form solution for the parameters that yield such a configuration or even the location of the light point. Nonetheless, by solving it numerically we found that for $a/M=1/2$ and $jM^{2}\simeq 0.158579$ the ergoregion develops precisely the type of saddle point described above, and the light point is indeed present there. This is presented in Fig. 7. To indicate the absence of angular velocity the light point is represented by an open circle.
Figure 7: Ergoregion of the KBHSU for $a/M=1/2$ and $jM^{2}\simeq 0.158579$, highlighted in blue. Here, the appearance of a saddle point is made clear, occurring when two previously disconnected regions merge. The light ring (blue circle) and light point (orange circle) of this solution are also shown. To highlight the absence of rotation the light point is represented by an open circle.
To the best of our knowledge, the appearance of light points had only been observed for spinning boson stars on the onset of the formation of an ergotorus [14], where the conditions $g_{tt}=0\>\wedge\>\nabla g_{tt}=0$ are also satisfied but correspond to a local maximum, instead of a saddle point. Light rings that appear at the onset of the ergoregion’s appearance are also stable, related to the fact that they correspond to local extrema of $g_{tt}$. Therefore, this represents the first instance in the literature of a light point in a black hole spacetime, which is unstable and possesses a distinct formation mechanism as compared to previous cases (merging of disconnected sections of ergoregions).
3.3 Topological charge
The fact that the light rings are critical points of some potential defined on the 2-dimensional space spanned by the non-Killing coordinates for the definition of a topological charge associated with the spacetime outside the possible event horizons. This topological charge yields strict bounds on the possible number of light rings on the spacetime.
For stationary, axially symmetric asymptotically flat spacetimes this topological charge has been used to establish that all spacetimes containing $N$ non-extremal event horizons always possess at least $N$ light rings per rotation sense [16, 21]. However, the asymptotic swirling behavior of the metric (1) places it outside the scope of those theorems. Nonetheless, such theorems have also been used to extract information on the number of light rings in other non-asymptotically flat, including asymptotically swirling spacetimes [20, 22, 23].
For the following analysis, it is more convenient to work with the metric in Bach-Weyl coordinates $\{\rho,z\}$[20]. The relation between the two coordinate systems is given by:
$\rho$ $=$ $\sqrt{\Delta}\sin\theta,$
$z$ $=$ $(r-M)\cos\theta,$ (22)
such that the metric now reads
${\rm d}s^{2}=\frac{1}{\mathcal{F}(r,\theta)}\left({\rm d}\varphi-\omega{\rm d}t\right)^{2}+\mathcal{F}(r,\theta)\left[-\rho^{2}{\rm d}t^{2}+\frac{\Sigma\sin^{2}\theta}{\Delta+(M^{2}-a^{2})\sin^{2}\theta}\left({\rm d}\rho^{2}+{\rm d}z^{2}\right)\right],$ (23)
where the $r$ and $\theta$ functions of $\rho$ and $z$ are implicitly given by the transformation in Eq. (22). The horizon corresponds now to the rod $\rho=0\,,-\sqrt{M^{2}-a^{2}}<z<\sqrt{M^{2}-a^{2}}$, while the remaining points with $\rho=0$ correspond to the rotation axis.
Let $\mathbf{v}^{\pm}$ be the vector field defined from the potential $H_{\pm}$ as:
$\mathbf{v}^{\pm}=\left(\frac{1}{\sqrt{g_{\rho\rho}}}\partial_{\rho}H_{\pm},\frac{1}{\sqrt{g_{zz}}}\partial_{z}H_{\pm}\right)=\left(v_{\rho}^{\pm},v_{z}^{\pm}\right)\,.$ (24)
This vector field can also be parametrized by its magnitude and an angle $\Omega$ such that:
$v^{\pm}_{\rho}=v^{\pm}\cos\Omega\,,\quad v^{\pm}_{z}=v^{\pm}\sin\Omega\,$, (25)
where $v^{\pm}=|\mathbf{v}^{\pm}|=\sqrt{(v^{\pm}_{\rho})^{2}+(v^{\pm}_{z})^{2}}$. Consider a closed curve, $\mathcal{C}$, in the $\{\rho,z\}$ plane. After a revolution along this curve, the angle $\Omega$ must be the same as in the beginning modulo $2\pi$, that is:
$\oint_{\mathcal{C}}\mathrm{d}\Omega=2\pi\omega_{\mathcal{C}}\,$, (26)
where $\omega\in\mathbb{Z}$ is the winding number of the vector field along the curve. The contributions to this winding number come from the critical points enclosed by $\mathcal{C}$, with an extremum (maximum or minimum) contributing with +1+1 and a saddle point contributing with 1-1. Notice that, since each critical point corresponds to the location of a light ring the winding number along some curve $\mathcal{C}$ can be used to extract information on the number of light rings it encloses.
In order to compute the total number of light rings on the spacetime outside the black hole, a contour $\mathcal{C}$ is considered such that in the appropriate limit, it encloses all the exterior spacetime. In Bach-Weyl coordinates, such a contour can be made of the union of the following six segments:
$\mathcal{I}_{1}=$ $\left\{\rho=M\xi\,,-M\xi<z<M\xi\right\}\,$, (27)
$\mathcal{I}_{2}=$ $\left\{\frac{M}{\xi}<\rho<M\xi\,,z=M\xi\right\}\,$, (28)
$\mathcal{I}_{3}=$ $\left\{\rho=\frac{M}{\xi}\,,\sqrt{M^{2}-a^{2}}<z<M\xi\right\}\,$, (29)
$\mathcal{I}_{4}=$ $\left\{\rho=\frac{M}{\xi}\,,-\sqrt{M^{2}-a^{2}}<z<\sqrt{M^{2}-a^{2}}\right\}\,$, (30)
$\mathcal{I}_{5}=$ $\left\{\rho=\frac{M}{\xi}\,,-M\xi<z<-\sqrt{M^{2}-a^{2}}\right\}\,$, (31)
$\mathcal{I}_{6}=$ $\left\{\frac{M}{\xi}<\rho<M\xi\,z=-\xi\right\}\,.$ (32)
That is $\mathcal{C}=\bigcup_{i}\mathcal{I}_{i}$. Notice that $\xi\in\mathbb{R}$ is a dimensionless parameter. In the limit $\xi\rightarrow\infty$, the contour $\mathcal{C}$ encloses all the spacetime outside the horizon, such that $\mathcal{I}_{1},\,\mathcal{I}_{2}$ and $\mathcal{I}_{6}$ capture the behavior of the potential at spatial infinity, $\mathcal{I}_{4}$ the horizon contribution, and the behavior near the rotational axis is captured by $\mathcal{I}_{3}$ together with $\mathcal{I}_{5}$. Under this decomposition, the total topological charge can also be decomposed in the contribution along each segment:
$w_{\mathcal{C}}=\oint_{\mathcal{C}}\mathrm{d}\Omega=\sum_{i}\omega_{\mathcal{I}_{i}}\,$, (33)
where
$\omega_{\mathcal{I}_{1}}$ $=\frac{1}{2\pi}\intop_{-M\xi}^{M\xi}\left.\frac{\mathrm{d}\Omega}{dz}\right|_{\rho=M\xi}\mathrm{dz}\>,$ $\omega_{\mathcal{I}_{2}}$ $=\frac{1}{2\pi}\intop_{M\xi}^{M/\xi}\left.\frac{\mathrm{d}\Omega}{d\rho}\right|_{z=M\xi}\mathrm{d\rho}\>,$
$\omega_{\mathcal{I}_{3}}$ $=\frac{1}{2\pi}\intop_{M\xi}^{\sqrt{M^{2}-a^{2}}}\left.\frac{\mathrm{d}\Omega}{dz}\right|_{\rho=M/\xi}\mathrm{dz}\>,$ $\omega_{\mathcal{I}_{4}}$ $=\frac{1}{2\pi}\intop_{\sqrt{M^{2}-a^{2}}}^{-\sqrt{M^{2}-a^{2}}}\left.\frac{\mathrm{d}\Omega}{dz}\right|_{\rho=M/\xi}\mathrm{dz}\>,\quad$ (34)
$\omega_{\mathcal{I}_{5}}$ $=\frac{1}{2\pi}\intop_{-\sqrt{M^{2}-a^{2}}}^{-M\xi}\left.\frac{\mathrm{d}\Omega}{dz}\right|_{\rho=M/\xi}\mathrm{dz}\>,$ $\omega_{\mathcal{I}_{6}}$ $=\frac{1}{2\pi}\intop_{M/\xi}^{M\xi}\left.\frac{\mathrm{d}\Omega}{d\rho}\right|_{z=-M\xi}\mathrm{d\rho}.$
Now one can study the behavior of the vector field $\mathbf{v^{\pm}}$ along each of the segments by making the appropriate perturbative expansions. Before continuing this analysis, it is important to notice that the effective potentials can be transformed into one another via the transformation $j\rightarrow-j$ and $\theta\rightarrow\pi-\theta$ (or equivalently $z\rightarrow-z$ in Bach-Weyl coordinates), i.e. $H_{+}^{a}(r,\theta)=-H_{-}^{-a}(r,\pi-\theta)$. Therefore, it is sufficient to analyze the properties of one of the potentials, for instance $H_{+}$. Moreover, the behavior of the vector field close to $\rho=0$ is strictly fixed by the regularity of the axis and the non-extremality of the horizon (which will be assumed from now on). Such a study has been done in [21]. There, it was concluded that each timelike rod (topologically spherical horizon) contributes with 1-1 to the total topological charge, so:
$\lim_{\xi\rightarrow\infty}\left(\omega_{\mathcal{I}_{3}}+\omega_{\mathcal{I}_{4}}+\omega_{\mathcal{I}_{5}}\right)=-1\,.$ (35)
Now we consider the asymptotic behavior of the vector field on the remaining segments:
Let us consider the behavior along $\mathcal{I}_{1}$, for $\rho=M\xi\rightarrow\infty$ one obtains:
$\left(v^{+}_{\rho},v^{+}_{z}\right)\sim\left(3|j|,0\right)\,.$ (36)
Since $\mathbf{v^{+}}$ is constant up to zeroth order in $\mathcal{O}(\xi^{-1})$ there is no contribution to the winding number along this segment, i.e. $\omega_{\mathcal{I}_{1}}=0$.
Next, we consider the segment $\mathcal{I}_{2}$, where $z\rightarrow\infty$, then:
$\left(v^{+}_{\rho},v^{+}_{z}\right)\sim\frac{1}{\sqrt{\left(1-4ajM\right)^{2}+\rho^{4}j^{2}}}\left(3j^{2}\rho^{2}-\frac{\left(1-4ajM\right)^{2}}{\rho^{2}},4j\left(1-4ajM\right)\right)\,.$ (37)
Notice that if $v^{\pm}_{z}>0\Rightarrow\Omega\in]0,\pi[$ and $v^{\pm}_{z}<0\Rightarrow\Omega\in]-\pi,0[$, then, it is possible to express the angle $\Omega\left(\rho,z\right)$ depending on the two possibilities for the value of 14ajM1-4ajM:
$\Omega\left(\rho,z\right)=\mathrm{sign}(j)\mathrm{sign}\left(1-4ajM\right)\arccos\left(\frac{v^{\pm}_{\rho}}{|\mathbf{v}^{\pm}|}\right)\,.$ (38)
Then one can compute $\omega_{\mathcal{I}_{2}}$ to be:
$2\pi\omega_{\mathcal{{I}_{2}}}=\left[\arccos\left(\frac{v^{+}_{\rho}}{|\mathbf{v}^{+}|}\right)\right]_{\rho=M\xi}^{\rho=M/\xi}=\mathrm{sign}(j)\mathrm{sign}\left(1-4ajM\right)\pi+\mathcal{O}\left(\xi^{-1}\right)\,.$ (39)
Only the behavior along $\mathcal{I}_{6}$ remains to be considered, this corresponds to the limit $z\rightarrow$-$\infty$, where:
$\left(v^{+}_{\rho},v^{+}_{z}\right)\sim\frac{1}{\sqrt{\left(1+4ajM\right)^{2}+\rho^{4}j^{2}}}\left(3j^{2}\rho^{2}-\frac{\left(1+4ajM\right)^{2}}{\rho^{2}},4j\left(1+4ajM\right)\right)\,.$ (40)
By the same procedure as before one finds:
$2\pi\omega_{\mathcal{I}_{6}}=\left[\arccos\left(\frac{v^{+}_{\rho}}{|\mathbf{v}^{+}|}\right)\right]^{\rho=M\xi}_{\rho=M/\xi}=-\mathrm{sign}(j)\mathrm{sign}\left(1+4ajM\right)\pi+\mathcal{O}\left(\xi^{-1}\right)\,.$ (41)
It is useful to compute the contribution from the three segments $\mathcal{I}_{1}\,,\mathcal{I}_{2}$ and $\mathcal{I}_{6}$ together, which yield:
$\lim_{\xi\rightarrow\infty}\left(\omega_{\mathcal{I}_{1}}+\omega_{\mathcal{I}_{2}}+\omega_{\mathcal{I}_{6}}\right)=\frac{\mathrm{sign}\left(j\right)}{2}\left[\mathrm{sign}\left(1-4ajM\right)-\mathrm{sign}\left(1+4ajM\right)\right]\,.$ (42)
There are three possibilities for the term in square brackets:
$\mathrm{sign}\left(1-4ajM\right)-\mathrm{sign}\left(1+4ajM\right)=\begin{cases}\ \ 2\,,\quad ajM<-1/4\\ \ \ 0\,,\quad-1/4<ajM<1/4\\ -2\,,\quad ajM>1/4\end{cases}\,.$ (43)
So one has three possibilities for the total topological charge of the spacetime depending on the value of the parameter $ajM$, which, considering $j>0$ are:
$\omega^{+}=\lim_{\xi\rightarrow\infty}\omega^{+}_{C}=\begin{cases}\ \ 0\,,\quad ajM<-1/4\\ -1\,,\quad-1/4<ajM<1/4\\ -2\,,\quad ajM>1/4\end{cases}\,.$ (44)
For $j<0$ the possibilities are the same but correspond to different regions of the parameter space.
At first glance this result seems to contradict the expectation that the topological charge is conserved under deformations of the spacetime, as one can change only one parameter and obtain three different values for the total topological charge for each potential. This occurs because the deformation is not smooth, in fact at the limiting values of the regions, i.e. $4ajM=\pm 1$ the metric is singular, as discussed previously. Another remark is that the light rings associated with the potential $H_{-}$ can be obtained by the substitution $a\rightarrow-a$ and $\theta\rightarrow\pi-\theta$, thus the topological charge associated with $\mathbf{v}^{-}$ is:
$\omega^{-}=\lim_{\xi\rightarrow 0}\omega^{+}_{C}=\begin{cases}-2\,,\quad ajM<-1/4\\ -1\,,\quad-1/4<ajM<1/4\\ \ \ 0\,,\quad ajM>1/4\end{cases}\,.$ (45)
Interestingly, the sum of the topological charges is independent of the value of $4ajM$, $\omega^{+}+\omega^{-}=-2$. The topological charge gives only the surplus of unstable light rings with respect to the stable ones, however, computing explicitly the location of the LRs for the considered spacetime, one never finds stable LRs. So there are only unstable LRs whose number is the modulus of the topological charge, which in this case is always $2$, despite the fact that the contribution from each potential may vary. In Fig. 8 the vector potentials are plotted for different values of $ajM$ such that all possibilities are covered.
Figure 8: Vector fields $\mathbf{v}^{\pm}$ for $jM^{2}=1$ and several values of the black hole spin parameter. On the left (right) column the vector $\mathbf{v}^{-}$ ($\mathbf{v}^{+}$) is plotted, with $a/M=\{-\sqrt{2}/2\,,\,-1/9,\,\sqrt{2}/2\}$ from top to bottom. It is possible to verify that all configurations possess in total 2 light rings, but for the top and bottom values of $a$ one potential contributes with both light rings, while for the middle row each potential contributes with one.
@@ 5 Conclusions
The spacetime of a Kerr black hole in a swirling universe is a closed-form solution in General Relativity, obtained by well-known transformations from the Kerr black hole seed solution [4]. Unlike the Schwarzschild black holes in a swirling universe, however, the interaction between the black hole spin and the swirling parameter does not yield a an (odd) $\mathbb{Z}_{2}$ symmetry of the spacetime. Thus the upper and lower hemispheres are not related by a simple symmetry transformation.
@@
Here we have investigated some properties of this spacetime. First, we have recalled the conical singularities, which diverge at a critical value of the modulus of the parameter combination $ajM=0.25$, where both spin parameters enter. We have computed the Kretschmann scalar and demonstrated that this conical singularity is a physical singularity. Next we have addressed the ergoregions. These also exhibit a notable change in the vicinity of the critical value. Whereas for small $ajM$, upper and lower hemispheres possess disconnected counter-rotating non-compact ergoregions in addition to the common ergosphere of the Kerr black hole, this changes at the critical value $ajM=0.25$. For larger values of $ajM$, the three disconnected regions are merged to become one infinitely extended ergoregion with hence only one sense of direction of rotation.
We have extensively studied the light rings in the KBHSU spacetimes. For the SBHSU spacetime the degenerate Schwarzschild LRs split up in a symmetrical way with increasing swirling parameter and move towards the horizon and the poles. For the KBHSU, the symmetry is broken and the two LRs now possess different radii and no longer have angular variables $\theta$ and $\pi-\theta$. Using a topological argument, we have demonstrated that two light rings should be present and that both are unstable. Furthermore, an interesting feature appears as the value of the swirling parameter is increased: there is a light point. This is a special regime in which a light ring exists exactly as the two sections of ergoregions merge; this LR does not possess a rotational sense. To the author’s knowledge, this is the first report of such a class of orbit.
@@
Finally, we investigated the shadow of KBHSU spacetimes. Since the solution is Petrov type I and hence the equations describing geodesic motion are not separable, we had to employ fully numerical methods to obtain the shadows seen by observers outside the black hole. The SBHSU shadows exhibit an odd-$\mathbb{Z}_{2}$ symmetry for observers in the equatorial plane. The interaction between the spins of the KBHSU spacetime does not allow for such odd-$Z_{2}$ symmetric shadows. All shadows are twisted, though, which would make the appearance of such a black hole most unusual.
@@
While the spacetime may not appear very realistic given the well-known features of our universe, it may still have features that might be of relevance. Here we would like to point out that the universe is threaded by huge structures, cosmic filaments, that rotate and thus might be modeled in part by the swirling universe solutions. Moreover, Astorino et al. [4] considered the swirling universe solutions as possibly relevant in the description of the collision of rotating galaxies. Finally, we would like to mention that it has recently been proposed that a rotating universe might solve the Hubble tension [32].
@@
The KBHSU is also interesting from another point of view : it is an example of a spacetime in which test particle motion is chaotic [33]. It would hence be an ideal system to quantify and qualify chaotic behaviour. This could lead to a better understanding of the underlying connection between chaotic test particle motion and the algebraic classification of solutions generated via suitable transformations in the Ernst formalism.
@@
Acknowledgments: R.C. would like to thank CAPES for financial support under Grant No: 88887.371717/2019-00. J.N. is supported by the FCT grant 10.54499/2021.06539.BD (https://doi.org/10.54499/2021.06539.BD).